OpenStudy (anonymous):

Help please! Solve: 50^2 + 60^2 = N times the square root of 625. A. 100 B. 120 C. 186 D. 244

OpenStudy (nurali):

50^2+60^2 = N*sqrt(625) 50^2=2500 60^2=3600 sqrt(625)=25 so plug and chug 2500+3600=N*25 6100/25=N N=244

OpenStudy (anonymous):

Thank you!

OpenStudy (nurali):

Anytime.